OpenStudy (anonymous):

Algebraically find the absolute maximum and minimum of f(x)=x-3ln(x) over [1,4]

OpenStudy (slaaibak):

Derive and set equal to zero: 0 = 1 - 3/x x = 3 Checking whether it's a min or max by calculating second derivative: f''(x) = 3/x^2 > 0 therefore it's a minimum. Subbing in the equation: f(3) = 3 - 3ln(3) = -0.295 Now checking the endpoints: f(1) = 1- 3ln1 = 1 f(4) = 4 - 3ln4 = -.158 So abs min: (3, -0.295) max: (1,1)